OpenStudy (toxicsugar22):

According to the following reaction, how many grams of water are necessary to form 0.478 moles oxygen gas? 2 H2O (l) = 2 H2 (g) + O2 (g)

OpenStudy (anonymous):

The above rxn tells us that " two molecules of water gives 2 molecules of hydrogen and 1 molecule of oxygen" It also says that " 2 moles of water gives 2 moles of hydrogen and 1 mole of oxygen" So 1 mole of O2 requires 2 moles of H2O Therefore 0.478 moles of O2 requires 2*0.478 moles of H2O =0.956 moles of H2O 1 mole of H2O = 18 grams 0.956 moles of H2O = 18 * 0.956 grams

OpenStudy (anonymous):

Here you do not have to worry about H2S as it is in excess. The quantity of water will be controlled by the amount of O2. The excess H2S will remain unreacted. Now proceed systematically as done before. 3 moles O2 gives 2 moles H2O OR 96 grams O2 gives 36 grams H2O 21.6 grams O2 will give (36*21.6)/96 grams of H2O

OpenStudy (anonymous):

Yes , i think so

OpenStudy (anonymous):

i don't know about any such calculator. Are u in eleventh standard ?

OpenStudy (anonymous):

Preparing for IIT ?

OpenStudy (anonymous):

Sure ,if i can do it. I am not too expert in chemistry.

OpenStudy (toxicsugar22):

According to the following reaction, how many grams of carbon dioxide will be formed upon the complete reaction of 24.3 grams of oxygen gas with excess carbon (graphite)? C (s) + O2 (g) = CO2 (g)

OpenStudy (anonymous):

Exactly like the previous problem. Do not worry about carbon as it is in excess. The dominant factor is the amount of O2. The quantity of CO2 will be controlled by amount of O2 taken. Can you do the rest of the part?

OpenStudy (toxicsugar22):

i am not sur if this is right as I have trouble with this kind of problem

OpenStudy (toxicsugar22):

is it 3 moles of oxygen gives 2 moles of carbon

OpenStudy (toxicsugar22):

or 96 grams of oxygen gives 36 grams carb on

OpenStudy (anonymous):

1 mole of O2 gives 1 mole of CO2 32 grams of O2 gives 44 grams of CO2

OpenStudy (anonymous):

There is no role of carbon. We are given some amount of O2 and we have to find the amount of CO2 it will give.

OpenStudy (toxicsugar22):

I got answer 33.41 is that right

OpenStudy (anonymous):

Absolutely.

OpenStudy (toxicsugar22):

yeah thank you sooo much for the help

OpenStudy (anonymous):

u are welcome !

OpenStudy (toxicsugar22):

what is IIT

OpenStudy (toxicsugar22):

YOU METIONED THAT